Solve for a and b :
-5b - 2a = 46
7a - 5b = 19

Found 2 solutions by mananth, stanbon:
Answer by mananth(12270)

-5 b -2 a = 46 .............1
-5 b + 7 a = 19 .............2
multiply equation (1) by -1
we get
5b+2a=-46
add to (2)
9a=-27
/9
a=-3
plug the value of a in (1)
-5b-2(-3)=46
-5b+6=46
-5b=40
/-5
b=-8

Answer by stanbon(57219)

Solve for a and b :
-5b - 2a = 46
7a - 5b = 19
---------------------
Rearrange:
-2a - 5b = 46
7a - 5b = 19
---------------------
Subtract and solve for "a":
9a = -27
a = -3
--------
Solve for "b":
-2a - 5b = 46
-2*-3 -5b = 46
6 - 5b = 46
-5b = 40
b = -8
------------------
Solution: (-3,-8)
=========================
Cheers,
Stan H.
=========================